40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Dec 19 19:51:19 2018, gautam, Update, General, Pumpdown prep OMCscan.pdf
    Reply  Wed Dec 19 22:11:28 2018, Koji, Update, General, How to align the copper OMC 110804_40m_OMC_layout.pdfOMC_alignment.pdf
Message ID: 14369     Entry time: Wed Dec 19 19:51:19 2018     Reply to this: 14371
Author: gautam 
Type: Update 
Category: General 
Subject: Pumpdown prep 

[Koji, gautam]

Summary:

We are ready to put the heavy doors back on the chambers and do some test pumpdowns tomorrow morning if Jon gives us the go-ahead. Also, Koji made the OMC resonate some of the AUX beam light we send into ityes

Details:

  1. EY work:
    • IMC was locked, and we attempted to locate the beam with an IR card inside the chamber.
    • Koji found that the beam was too high, we were over-shooting the entire black-glass baffle on the EY table.
    • So I moved the TTs to try and center the beam through the aperture of aforementioned baffle.
    • Once this was done, we found that the beam was misaligned in yaw by ~1-inch in transmission on the EY optics table (there was an iris in place marking the cavity transmission axis). This explains why I couldn't find any TRY flashes while moving the TTs around.
    • We hypothesize that without the 2 degree ETM wedge in place, there isn't a compatible axis for the ITM transmission to also make it through the EY baffle and transmission iris. Over ~1m, the 2 degree wedge makes roughly 1.4 inch translation in yaw, so this seems to be a plausible hypothesis.
    • The ETMY suspension was moved from the mini-cleanroom setup back into the EY vacuum chamber. Two clamps (finger tightened only) hold it in place on the NE edge of the optical table. We decided that this is a better resting palce for the cage over the holidays than an in-air cleanroom.
  2. OMC chamber work:
    • While we were in clean garb, we decided to also investigate the OMC situation a bit.
    • It quickly became apparent that it was hopeless for me to work in chamber in the tightly confined IOO chamber. So Koji went in to have a look.
    • Koji will post the detailed alignment procedure - but after some alignment of the AUX laser input beam axis using in air steering mirrors and Koji's expert tweaking of the pointing into the OMC, we observed some resonances of the OMC.
    • Attachment #1 shows the full-range triangle ramp applied to the OMC length PZT (top row) and the OMC REFL signal (bottom row), measured using a PDA520 (chosen for its large active area) connected to a scope (AC-coupled, 1Mohm impedance, averaged to make the dips more prominent).
    • The OMC transmission was also (barely) visible on an IR card.
    • So the OMC length PZT seems capable of sweeping the length of the cavity. Based on the size of the dips we saw, the MM into the cavity is sub 1-percent.
    • The transmission PDs didn't output any measurable signal - but I'm not sure that the satellite box / readout electronics have been carefully characterized on the electroncis bench, so that will have to be done first.
    • We replaced the copper cover of the OMC (finger tightened for now) in case we do any test pumpdowns tomorrow. HV supply has been turned off, and the AUX laser has been reverted to standby mode.
Attachment 1: OMCscan.pdf  158 kB  | Hide | Hide all
OMCscan.pdf
ELOG V3.1.3-